OpenStudy (anonymous):

Decompose into partial fractions. (5x^3 -x^2 + 8x -55) / (x^4 + 5x^3 + 11x^2)

OpenStudy (anonymous):

\[\frac{2 x-11}{x^2+5 x+11}-\frac{5}{x^2}+\frac{3}{x} \]

OpenStudy (dumbcow):

\[\frac{5x^{3}-x^{2}+8x-55}{x^{4} +5x^{3}+11x^{2}} = \frac{5x^{3}-x^{2}+8x-55}{x^{2}(x^{2}+5x+11)} = \frac{A}{x^{2}}+\frac{B}{x}+\frac{Cx+D}{x^{2}+5x+11}\]

OpenStudy (anonymous):

Ok so what do i do after that?

OpenStudy (anonymous):

I actually dont even know what decmpose into partial fractions mean...i assumed it meant breaking it up to smaller fractions but i'm not sure how to do that. Is taht the end of the problem?

OpenStudy (dumbcow):

\[= \frac{A(x^{2}+5x+11) + Bx(x^{2}+5x+11)+(Cx+D)x^{2}}{x^{2}(x^{2}+5x+11}\]

OpenStudy (dumbcow):

yeah its breaking it into fractions which are equivalent to the original single fraction

OpenStudy (dumbcow):

Now find values for A,B,C,D so it matches original numerator \[A(x^{2}+5x+11) + Bx(x^{2}+5x+11)+(Cx+D)x^{2} = 5x^{3} -x^{2}+8x-55\]

OpenStudy (dumbcow):

\[(B+C)x^{3}+(A+5B+D)x^{2}+(5A+11B)x +11A = 5x^{3}-x^{2}+8x-55\] \[B+C = 5\] \[A+5B+D = -1\] \[5A+11B = 8\] \[11A = -55\]

OpenStudy (dumbcow):

A = -5 B = 3 C=2 D = -11

OpenStudy (dumbcow):

\[= -\frac{5}{x^{2}}+\frac{3}{x}+\frac{2x-11}{x^{2}+5x+11}\]

OpenStudy (anonymous):

THANK YOU SOOO MUCH! I'm a visual learner so this was such a great step by step process. I really appreciate your help.

OpenStudy (dumbcow):

your welcome

OpenStudy (anonymous):

I have another question..do u mind? It's just to confirm ansnwer about the domain and range of a function.

OpenStudy (dumbcow):

yeah sure

OpenStudy (anonymous):

Is the domain and range of a function any number that makes it positive if the function is a sqrt root. For example the sqrt root of x^2 +44, \[dom: x \ge \sqrt{44} and range is y \ge 0\]

OpenStudy (dumbcow):

ok range is correct, range of sqrt is all positive numbers domain is off, since x^2 +44 is always greater than zero, x is all real numbers

OpenStudy (anonymous):

but doesnt real #'s include neg. #'s and a sqt can't be negative

OpenStudy (dumbcow):

that is true, but here it applies to x^2+44 not just x x could be anything

OpenStudy (anonymous):

oh i see now... so if it were like \[f(x) = \sqrt{x ^{2}-44}\] then domain would be x\[x \ge -2\]

OpenStudy (dumbcow):

i am wrong about the range, it should be y>= sqrt(44) it helps if you can graph the function to see domain and range

OpenStudy (anonymous):

is my previous note correct with teh domain?

OpenStudy (dumbcow):

not quite x^2 -44 > 0 x^2 > 44 x > sqrt(44) and x< -sqrt(44)

OpenStudy (dumbcow):

hope this helps got to go

OpenStudy (anonymous):

THANKS ALOT ! =) IT REALLY DID